# P-ICL: Point In-Context Learning for Named Entity Recognition with Large Language Models

### v1

@@ 6 Conclusion
In this paper, we introduce P-ICL, a framework that employs point entities to improve LLM in performing NER through ICL. By incorporating point entities, we equip LLMs with specific entity type information and entity classification, addressing the limitations of the standard ICL approach for NER. Our findings, based on four widely used NER datasets and experiments conducted with three renowned LLMs, confirm the efficacy of P-ICL.
@@
Appendix A Overall Performance
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | Avg. | |
Vanilla | 62.49 | 8.70 | 17.99 | 14.35 | 25.88 | |
Standard ICL | 3-shot | 64.95 | 10.55 | 20.30 | 19.84 | 28.91 |
5-shot | 66.47 | 10.92 | 22.48 | 22.32 | 30.55 | |
10-shot | 67.65 | 11.51 | 23.94 | 25.42 | 32.13 | |
20-shot | 66.35 | 10.39 | 23.90 | 27.05 | 31.92 | |
P-ICL | 5+3-shot | 66.80 | 16.61 | 24.31 | 24.57 | 33.07 |
10+3-shot | 65.70 | 14.93 | 23.59 | 24.19 | 32.10 | |
5+5-shot | 67.56 | 16.95 | 25.49 | 27.01 | 34.25 | |
10+5-shot | 66.58 | 16.10 | 25.85 | 26.51 | 33.76 | |
5+10-shot | 67.66 | 16.83 | 28.18 | 28.95 | 35.40 | |
10+10-shot | 68.27 | 18.51 | 26.64 | 30.02 | 35.86 | |
5+20-shot | 69.31 | 19.36 | 29.10 | 32.05 | 37.46 | |
10+20-shot | 69.58 | 18.40 | 28.31 | 31.99 | 37.07 | |
P-ICL | 5+3-shot | 68.87 | 19.10 | 26.34 | 25.57 | 34.97 |
10+3-shot | 69.47 | 17.99 | 26.88 | 25.36 | 34.93 | |
5+5-shot | 70.33 | 16.70 | 27.23 | 29.12 | 35.85 | |
10+5-shot | 70.83 | 19.41 | 26.74 | 28.09 | 36.27 | |
5+10-shot | 71.04 | 18.18 | 29.66 | 31.03 | 37.48 | |
10+10-shot | 71.12 | 20.98 | 29.11 | 30.45 | 37.92 | |
5+20-shot | 72.58 | 21.09 | 31.17 | 32.69 | 39.38 | |
10+20-shot | 71.80 | 20.23 | 30.53 | 32.19 | 38.69 |
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | Avg. | |
Vanilla | 32.12 | 6.83 | 22.07 | 18.77 | 19.95 | |
Standard ICL | 3-shot | 43.65 | 6.34 | 31.17 | 27.97 | 27.28 |
5-shot | 46.54 | 6.95 | 34.25 | 30.95 | 29.67 | |
10-shot | 50.88 | 8.02 | 37.86 | 36.43 | 33.30 | |
20-shot | 50.13 | 7.88 | 37.57 | 36.18 | 32.94 | |
P-ICL | 5+3-shot | 47.10 | 7.96 | 36.10 | 32.25 | 30.85 |
10+3-shot | 46.94 | 7.51 | 34.81 | 32.68 | 30.49 | |
5+5-shot | 50.41 | 8.94 | 37.81 | 34.84 | 33.00 | |
10+5-shot | 49.31 | 7.51 | 35.69 | 35.89 | 32.10 | |
5+10-shot | 54.32 | 8.62 | 39.69 | 36.05 | 34.67 | |
10+10-shot | 54.92 | 7.77 | 39.03 | 38.15 | 34.97 | |
5+20-shot | 58.63 | 8.53 | 40.44 | 39.66 | 36.82 | |
10+20-shot | 58.44 | 8.29 | 40.35 | 39.60 | 36.67 | |
P-ICL | 5+3-shot | 49.96 | 9.45 | 36.04 | 32.45 | 31.97 |
10+3-shot | 49.02 | 8.83 | 36.88 | 32.59 | 31.83 | |
5+5-shot | 52.11 | 9.92 | 39.68 | 35.54 | 34.31 | |
10+5-shot | 51.37 | 9.92 | 39.38 | 36.46 | 34.28 | |
5+10-shot | 57.62 | 9.35 | 41.05 | 38.14 | 36.54 | |
10+10-shot | 56.41 | 10.30 | 40.98 | 39.12 | 36.70 | |
5+20-shot | 61.70 | 8.40 | 42.32 | 40.67 | 38.27 | |
10+20-shot | 60.40 | 10.74 | 42.51 | 40.99 | 38.66 |

### v2

@@ 6 Conclusion
In this paper, we introduce P-ICL, a framework that employs point entities to improve LLM in performing NER through ICL. By incorporating point entities, we equip LLMs with specific entity type information and entity classification, addressing the limitations of the standard ICL approach for NER. Our findings, based on six widely used NER datasets and experiments conducted with three renowned LLMs, confirm the efficacy of P-ICL.
@@
Appendix A Implementation Details
We report the entity-level micro F1 scores in the following result tables and figures to compare the NER performance. To more comprehensively compare the P-ICL performance of existing LLMs, we selected three representative models at this stage, including GPT-3.5 Turbo, LLaMA-3-70B AI@Meta (2024) and Mixtral 8x7B Jiang et al. (2024a). The demonstrations used in ICL are selected randomly in the training set. For standard ICL and P-ICL experiments, due to a certain degree of randomness, the corresponding experimental results are the average of the F1 scores of 5 experiments. For experiments on open-source models, we use the vLLM Kwon et al. (2023) framework for inference. In order to maintain the stability and consistency of model generation, we use the greedy generation to control the outputs of LLMs. We conducted our experiments on eight NVIDIA A800 80GB PCIe GPUs.
Appendix B Overall Performance
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | JNLPBA | BC5CDR | Avg. | |
Vanilla | 62.49 | 8.70 | 17.99 | 14.35 | 38.83 | 68.30 | 35.11 | |
Standard ICL | 3-shot | 64.95 | 10.55 | 20.30 | 19.84 | 42.88 | 67.47 | 37.67 |
5-shot | 66.47 | 10.92 | 22.48 | 22.32 | 43.53 | 68.17 | 38.98 | |
10-shot | 67.65 | 11.51 | 23.94 | 25.42 | 46.09 | 68.67 | 40.55 | |
20-shot | 66.35 | 10.39 | 23.90 | 27.05 | 45.05 | 69.12 | 40.31 | |
P-ICL | 5+3-shot | 66.80 | 16.61 | 24.31 | 24.57 | 45.55 | 67.10 | 40.82 |
10+3-shot | 65.70 | 14.93 | 23.59 | 24.19 | 45.38 | 66.41 | 40.03 | |
5+5-shot | 67.56 | 16.95 | 25.49 | 27.01 | 46.83 | 67.64 | 41.91 | |
10+5-shot | 66.58 | 16.10 | 25.85 | 26.51 | 46.05 | 67.40 | 41.42 | |
5+10-shot | 67.66 | 16.83 | 28.18 | 28.95 | 48.00 | 69.10 | 43.12 | |
10+10-shot | 68.27 | 18.51 | 26.64 | 30.02 | 48.01 | 68.33 | 43.30 | |
5+20-shot | 69.31 | 19.36 | 29.10 | 32.05 | 48.99 | 69.60 | 44.73 | |
10+20-shot | 69.58 | 18.40 | 28.31 | 31.99 | 49.16 | 69.28 | 44.45 | |
P-ICL | 5+3-shot | 68.87 | 19.10 | 26.34 | 25.57 | 48.17 | 69.73 | 42.96 |
10+3-shot | 69.47 | 17.99 | 26.88 | 25.36 | 47.85 | 68.34 | 42.65 | |
5+5-shot | 70.33 | 16.70 | 27.23 | 29.12 | 48.39 | 70.42 | 43.70 | |
10+5-shot | 70.83 | 19.41 | 26.74 | 28.09 | 48.65 | 69.18 | 43.82 | |
5+10-shot | 71.04 | 18.18 | 29.66 | 31.03 | 50.22 | 71.27 | 45.23 | |
10+10-shot | 71.12 | 20.98 | 29.11 | 30.45 | 51.04 | 70.36 | 45.51 | |
5+20-shot | 72.58 | 21.09 | 31.17 | 32.69 | 51.06 | 71.26 | 46.64 | |
10+20-shot | 71.80 | 20.23 | 30.53 | 32.19 | 50.86 | 71.10 | 46.12 |
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | JNLPBA | BC5CDR | Avg. | |
Vanilla | 37.38 | 6.31 | 19.01 | 17.26 | 21.85 | 54.63 | 26.07 | |
Standard ICL | 3-shot | 45.93 | 8.76 | 28.22 | 25.42 | 26.95 | 57.69 | 32.16 |
5-shot | 48.10 | 10.11 | 30.54 | 27.90 | 28.16 | 57.68 | 33.75 | |
10-shot | 51.00 | 12.56 | 32.42 | 30.94 | 29.38 | 57.83 | 35.69 | |
20-shot | 51.12 | 11.37 | 32.93 | 29.78 | 32.08 | 56.38 | 35.61 | |
P-ICL | 5+3-shot | 46.37 | 10.62 | 31.09 | 28.79 | 25.89 | 59.11 | 33.65 |
10+3-shot | 45.85 | 11.00 | 30.74 | 28.91 | 25.46 | 56.91 | 33.15 | |
5+5-shot | 48.87 | 11.67 | 31.13 | 31.33 | 28.15 | 58.84 | 35.00 | |
10+5-shot | 48.53 | 12.64 | 33.26 | 31.06 | 27.88 | 58.30 | 35.28 | |
5+10-shot | 53.09 | 13.49 | 35.92 | 34.57 | 29.86 | 59.45 | 37.73 | |
10+10-shot | 53.07 | 13.09 | 35.37 | 35.04 | 31.33 | 57.96 | 37.64 | |
5+20-shot | 56.50 | 14.85 | 37.02 | 36.69 | 32.08 | 58.28 | 39.24 | |
10+20-shot | 56.67 | 14.26 | 36.01 | 35.22 | 30.44 | 57.99 | 38.43 | |
P-ICL | 5+3-shot | 50.59 | 12.54 | 32.18 | 31.27 | 27.87 | 62.14 | 36.10 |
10+3-shot | 48.59 | 13.00 | 33.67 | 30.69 | 29.38 | 58.17 | 35.58 | |
5+5-shot | 51.81 | 14.55 | 36.08 | 34.03 | 29.21 | 61.36 | 37.84 | |
10+5-shot | 50.02 | 13.75 | 37.12 | 33.59 | 31.89 | 59.23 | 37.60 | |
5+10-shot | 56.18 | 15.36 | 36.77 | 36.78 | 32.40 | 61.10 | 39.77 | |
10+10-shot | 54.42 | 15.06 | 37.82 | 36.14 | 32.99 | 59.79 | 39.37 | |
5+20-shot | 59.12 | 16.92 | 39.38 | 38.60 | 35.05 | 60.41 | 41.58 | |
10+20-shot | 58.70 | 16.19 | 41.18 | 38.34 | 32.26 | 59.18 | 40.98 |